1. You own a portfolio that has $1,700 invested in Stock A and $3,000 invested in...

1.

You own a portfolio that has $1,700 invested in Stock A and $3,000 invested in Stock B. If the expected returns on these stocks are 10 percent and 18 percent, respectively, what is the expected return on the portfolio?(Do not round your intermediate calculations.)


rev: 09_20_2012

15.11%

12.89%

14.00%

15.86%

15.41%

2.

Suppose a stock had an initial price of $55 per share, paid a dividend of $1.75 per share during the year, and had an ending share price of $71. Compute the percentage total return.


rev: 09_20_2012

25.00

40.73

32.27

33.89

3.

The Up and Coming Corporation's common stock has a beta of 1.5. If the risk-free rate is 3 percent and the expected return on the market is 11 percent, what is the company's cost of equity capital? (Do not round your intermediate calculations.)


rev: 09_20_2012

15.75%

14.25%

15.6%

19.5%

15%
Answer #1

1.Total investment=(1700+3000)=$4700

Hence expected return of portfolio=Respective expected return*Respective investment weight

=(1700/4700*10)+(3000/4700*18)

=15.11%(Approx).

2.% total return=(End price-Beginning price+Dividends)/Beginning price

=(71-55+1.75)/55

=32.27%(Approx).

3. cost of equity capital=risk-free rate +Beta*(market rate- risk-free rate )

=3+1.5*(11-3)

=15%